# Structural and rigidity properties of Lie skew braces

### v1

@@ 2. $(G,\cdot)$ ab, $(G,\circ)$ solv
for all $a,a^{\prime},b,b^{\prime}\in\mathbb{R}$. The group $(G,\circ)$ is solvable but not nilpotent. By Example 4, we deduce that this structure induces a LSB on $\mathbb{R}^{2}$, denoted $(\mathbb{R}^{2},+,\circ)$, where $+$ is the standard addition on $\mathbb{R}^{2}$ and $\circ$ is the nontrivial solvable affine product defined above. Similar examples can be constructed, via pre-Lie algebra techniques, whenever Milnor’s question has a positive answer.
@@ 3. $(G,\cdot)$ nil, $(G,\circ)$ ab
By applying Example 3 we obtain a LSB $(\mathbb{R}^{3},\,\cdot,\,+),$ in which $(\mathbb{R}^{3},\cdot)$ is nilpotent and non-abelian, whereas $(\mathbb{R}^{3},+)$ is abelian.
@@ 4. $(G,\cdot)$ nil, $(G,\circ)$ nil
In [5], it is proved that for every $n\leq 5$, and for any pair of connected and simply connected nilpotent Lie groups $(G,\circ_{G})$ and $(K,\cdot_{K})$ of dimension $n$, there exists a simply transitive affine action
@@ 5. $(G,\cdot)$ nil, $(G,\circ)$ solv
Let $(G,\circ)$ be any connected and simply connected solvable Lie group that is not nilpotent. Then, by applying statement (F2) of Theorem 1.3, one can construct a LSB $(G,\cdot,\circ)$ such that the group $(G,\cdot)$ is nilpotent.
@@ 6. $(G,\cdot)$ solv, $(G,\circ)$ ab
Consider the affine group $(G,\cdot)=\mathrm{Aff}(\mathbb{R},+)\cong(\mathbb{R},+)\rtimes_{\alpha}(\mathbb{R},+)$, i.e., the group of automorphisms of the additive group $(\mathbb{R},+)$, as in the previous example. This time, however, we denote the affine group operation by $\cdot$ to emphasize its role as the primary group structure. By applying Example 3, we deduce that this setting induces a LSB structure on $\mathbb{R}^{2}$, denoted $(\mathbb{R}^{2},\cdot,+)$, where $(\mathbb{R}^{2},\cdot)$ is a solvable Lie group and $(\mathbb{R}^{2},+)$ is abelian.
@@ 7. $(G,\cdot)$ solv, $(G,\circ)$ nil
By applying Example 3 we obtain a LSB $(\mathbb{R}^{4},\,\cdot,\,\circ),$ in which $(\mathbb{R}^{4},\cdot)$ is solvable and non-nilpotent, whereas $(\mathbb{R}^{4},\circ)$ is nilpotent and non-abelian.
@@ 8. $(G,\cdot)$ solv, $(G,\circ)$ solv
Let $H$ be a connected, solvable, but non-nilpotent Lie group. Define the inner automorphism action
$\operatorname{Ad}\colon H\longrightarrow\mathrm{Aut}(H),\qquad\operatorname{Ad}_{a}(b)=a\,b\,a^{-1},$
and form the associated semidirect product $(G,\cdot)\;:=\;H\rtimes_{\operatorname{Ad}}H$. Equip the same manifold $H\times H$ with the direct-product structure $(G,\circ)\;:=\;H\times H$. Both $(G,\circ)$ and $(G,\cdot)$ are connected and solvable. The action $\operatorname{Ad}$ is non-trivial because $H$ is not abelian and hence $(G,\circ)\not\cong(G,\cdot)$. Applying Example 3, we obtain a non-trivial LSB $(G,\cdot,\circ),$ in which $(G,\cdot)$ and $(G,\circ)$ are distinct yet both solvable Lie groups.
@@ 9. $(G,\cdot)$ simple, $(G,\circ)$ solv and $(G,\cdot)$ ssimple, $(G,\circ)$ solv
The subgroup $K\cong\mathbb{S}^{1}$ is compact, connected, and abelian, and therefore nilpotent. The product $AN$ forms a connected, simply-connected, solvable Lie subgroup. Moreover, $K\cap AN=\{I_{2}\}$, where $I_{2}$ denotes the $2\times 2$ identity matrix. Hence, Example 3 applies, yielding a LSB structure $\left(\mathrm{SL}_{2}(\mathbb{R}),\cdot,\circ\right),$ where $\left(\mathrm{SL}_{2}(\mathbb{R}),\cdot\right)$ is simple, while $\left(\mathrm{SL}_{2}(\mathbb{R}),\circ\right)\cong K\times AN$ is solvable. Finally, taking the product of the previous LSB with itself (cf. Example 2) yields $\bigl{(}\mathrm{SL}_{2}(\mathbb{R})\times\mathrm{SL}_{2}(\mathbb{R})\,\cdot,\,\circ\bigr{)},$ where $\bigl{(}\mathrm{SL}_{2}(\mathbb{R})\times\mathrm{SL}_{2}(\mathbb{R}),\,\cdot\bigr{)}$ is semisimple and $(\mathrm{SL}_{2}(\mathbb{R})\times\mathrm{SL}_{2}(\mathbb{R}),\,\circ\bigr{)}$ is solvable.
@@ 10. $(G,\cdot)$ simple, $(G,\circ)$ mixed-type and $(G,\cdot)$ ssimple, $(G,\circ)$ mixed-type
  • $N$ is a three-dimensional, step-two nilpotent subgroup (isomorphic to the Heisenberg group $H_{3}$).
@@
Hence Example 3 applies and yields a LSB structure $\bigl{(}\mathrm{SL}_{3}(\mathbb{R}),\,\cdot,\,\circ\bigr{)},$ ssuch that $\bigl{(}\mathrm{SL}_{3}(\mathbb{R}),\cdot\bigr{)}$ is simple, whereas $\bigl{(}\mathrm{SL}_{3}(\mathbb{R}),\circ\bigr{)}\;\cong\;K\times AN$ is of mixed type: it contains both a non-abelian semisimple factor ($K$) and a solvable factor ($AN$), so it is neither solvable, nor simple, nor semisimple. Analogously with the previous example by taking the product of $\bigl{(}\mathrm{SL}_{3}(\mathbb{R}),\,\cdot,\,\circ\bigr{)},$ with itself one gets a LSB such that $\bigl{(}\mathrm{SL}_{3}(\mathbb{R})\times\mathrm{SL}_{3}(\mathbb{R}),\,\cdot\bigr{)}$ is semisimple and $(\mathrm{SL}_{3}(\mathbb{R})\times\mathrm{SL}_{3}(\mathbb{R}),\,\circ\bigr{)}$ is of mixed-type.
@@ 12. $(G,\cdot)$ mixed-type, $(G,\circ)$ mixed-type
By applying Example 3 we obtain a LSB $(\mathbb{R}^{3}\times\mathrm{SU}(2),\cdot,\circ)$ where both $(\mathbb{R}^{3}\times\mathrm{SU}(2),\cdot)$ and $(\mathbb{R}^{3}\times\mathrm{SU}(2),\circ)$ are of mixed-type and not isomorphic.
[…]
Step 3. The non-existence cases ()({-})( - ). Every remaining entry marked by a dash ()({-})( - ) represents a non-existence phenomenon. Such combinations are ruled out by $\mathbf{(S2)}$ together with $\mathbf{(R1)}$ of Theorem 1.1.

### v3

@@ 2. $(G,\cdot)$ ab, $(G,\circ)$ solv
for all $a,a^{\prime},b,b^{\prime}\in\mathbb{R}$. The group $(G,\circ)$ is solvable but not nilpotent. By Example 4, we deduce that this structure induces an LSB on $\mathbb{R}^{2}$, denoted $(\mathbb{R}^{2},+,\circ)$, where ++ is the standard addition on $\mathbb{R}^{2}$ and $\circ$ is the nontrivial solvable affine product defined above. Similar examples can be constructed, via pre-Lie algebra techniques, whenever Milnor’s question has a positive answer.
@@ 3. $(G,\cdot)$ nil, $(G,\circ)$ ab
[…]
@@ 4. $(G,\cdot)$ nil, $(G,\circ)$ nil
In [5, Proposition 4.1], it is proved that for every $n\leq 5$, and for any pair of connected and simply connected nilpotent Lie groups $(G,\circ_{G})$ and $(K,\cdot_{K})$ of dimension $n$, there exists a simply transitive affine action
@@ 5. $(G,\cdot)$ nil, $(G,\circ)$ solv
Let $(G,\circ)$ be any connected and simply connected solvable Lie group that is not nilpotent. Then, by applying statement (F2) of Theorem 1.3, one can construct an LSB $(G,\cdot,\circ)$ such that the group $(G,\cdot)$ is nilpotent.
@@ 6. $(G,\cdot)$ solv, $(G,\circ)$ ab
Consider the affine group $(G,\cdot)=\mathrm{Aff}(\mathbb{R},+)\cong(\mathbb{R},+)\rtimes_{\alpha}(\mathbb{R},+)$, i.e., the group of automorphisms of the additive group $(\mathbb{R},+)$, as in item 2. above. This time, however, we denote the affine group operation by $\cdot$ to emphasize its role as the primary group structure. By applying Example 3, we deduce that this setting induces an LSB structure on $\mathbb{R}^{2}$, denoted $(\mathbb{R}^{2},\cdot,+)$, where $(\mathbb{R}^{2},\cdot)$ is a solvable Lie group and $(\mathbb{R}^{2},+)$ is abelian.
@@ 7. $(G,\cdot)$ solv, $(G,\circ)$ nil
[…]
@@ 8. $(G,\cdot)$ solv, $(G,\circ)$ solv
Following [27, Example 2.19], let $n\geq 3$ and let $G:=UT_{n}^{+}(\mathbb{R})$ be the Lie group of invertible upper triangular $n\times n$ real matrices with positive diagonal entries. Every $A\in G$ can be written uniquely as
$A=N(A)+D(A),$
where $N(A)$ is strictly upper triangular and $D(A)$ is diagonal. Let $\circ$ denote the usual matrix multiplication on $G$, and define a second operation $\cdot$ on $G$ by
$A\cdot B:=N(A)+BD(A)\qquad(A,B\in G).$
It is straightforward to check that $\cdot$ is a Lie group law on $G$, and that $(G,\cdot,\circ)$ is an LSB, where $(G,\circ)$ is the standard solvable (non-nilpotent) upper triangular Lie group. Moreover, one can show that the commutator in $(G,\cdot)$ is given by
$A,B]_{\cdot}=\mathrm{I}_{n}+N(B)\bigl(D(A)-\mathrm{I}_{n}\bigr)-N(A)\bigl(D(B)-\mathrm{I}_{n}\bigr).$
Thus, one can verify that $(G,\cdot)$ is solvable but not nilpotent. Finally, one easily checks that $(G,\cdot)$ is metabelian: indeed, its commutator subgroup is
$\{\,\mathrm{I}_{n}+N\mid N\text{ is strictly upper triangular}\,\},$
which is abelian with respect to $\cdot$. On the other hand, $(G,\circ)$ is not metabelian for $n\geq 3$, since its commutator subgroup is the unitriangular group, which is non-abelian. Hence $(G,\cdot)\not\simeq(G,\circ)$, and we obtain a non-trivial LSB $(G,\cdot,\circ)$ with both groups solvable and not nilpotent.
@@ 9. $(G,\cdot)$ simple, $(G,\circ)$ solv and $(G,\cdot)$ ssimple, $(G,\circ)$ solv
[…]
@@ 10. $(G,\cdot)$ simple, $(G,\circ)$ mixed-type and $(G,\cdot)$ ssimple, $(G,\circ)$ mixed-type
  • NN is a three-dimensional, step-two nilpotent subgroup (isomorphic to the Heisenberg group $H_{3}$).
@@
[…]
@@ 12. $(G,\cdot)$ mixed-type, $(G,\circ)$ mixed-type
By applying Example 3 we obtain an LSB $(\mathbb{R}^{3}\times\mathrm{SU}(2),\cdot,\circ)$ where both $(\mathbb{R}^{3}\times\mathrm{SU}(2),\cdot)$ and $(\mathbb{R}^{3}\times\mathrm{SU}(2),\circ)$ are of mixed-type and not isomorphic.
Step 2. The trivial cases $({\cong})$. A connected LSB $\bigl(G,\cdot,\circ\bigr)$ is trivial whenever the two underlying Lie group structures $\bigl(G,\cdot\bigr)$ and $\bigl(G,\circ\bigr)$ are both abelian. Indeed, on a connected manifold any two abelian Lie group laws are isomorphic. The only other trivial situations are those in which $\bigl(G,\cdot\bigr)$ and $\bigl(G,\circ\bigr)$ are simultaneously semisimple (or even simple). These follow directly from condition $\mathbf{(R2)}$ in Theorem 1.1.
Step 3. The non-existence cases ()({-}). Every remaining entry marked by a dash ()({-}) represents a non-existence phenomenon. Such combinations are ruled out by $\mathbf{(S2)}$ together with $\mathbf{(R1)}$ of Theorem 1.1.
Acknowledgements
We would like to thank Giovanni Placini and Fabio Zuddas for carefully reading our manuscript and for their valuable comments and suggestions. We also thank the anonymous referee for a thorough reading and for helpful remarks that improved the exposition of the paper. In particular, we are grateful to the referee for suggesting the inclusion of item 8. in the proof of Theorem 1.4.